OpenStudy (anonymous):

dy/dx=(y-1)^2 cos(piex) Find the particular solution y=f(x) wo the differential equation with the initial condition f(1)=0.

OpenStudy (anonymous):

to*

OpenStudy (anonymous):

Oh hi, there's your question! dy = (y-1)^2 cos(pi * x) dx dy / (y-1)^2 = cos(pi *x) dx -(y-1)^-1 = sin(pi * x) +C Plug in x = 1 and y = 0 to find C -(-1)^-1 = sin(pi) + C 1 = 0 + c C = 1

OpenStudy (anonymous):

Become a fan :D

OpenStudy (anonymous):

hi! thanks. how do i know when to solve for c? and i will. your awesome!